Find the width of a rectangle with a length of 6yd and a perimeter of 18yds
Answer by ewatrrr(24785)

 
Hi,
.
P =2L + 2w = 18yd
L = 6yd
...
12 + 2w = 18
2w = 6
w = 3yd
and...checking
2L + 2w = P
12yd + 6yd = 18yd